OpenStudy (anonymous):

martin has 3 pounds of pasta 1.9 of cheese and 10 ounces of spice mix if he cooks everything and mixes it all together about h0w many pounds of spicy macaroni and cheese will he have?

hero (hero):

That's disgusting...mixing spaghetti and macaroni together.

OpenStudy (anonymous):

ikr but its a question though

hero (hero):

I say convert everything to ounces first, then add, then convert back to pounds

hero (hero):

pasta = 48 ounces cheese = 30.4 ounces sauce = 10 ounces 48 + 30.4 + 10 = 58 + 30.4 = 88.4 88.4/16 = 5.525 So a total of 5.525 pounds of the nasty stuff

OpenStudy (anonymous):

wrong answer